# Griesmer and Optimal Linear Codes from the Affine Solomon-Stiffler Construction

### v2

@@ Abstract
In their fundamental paper published in 1965, G. Solomon and J. J. Stiffler invented infinite families of codes meeting the Griesmer bound. These codes are then called Solomon-Stiffler codes and have motivated various constructions of codes meeting or close the Griesmer bound.
In this paper, we give a geometric construction of infinite families of affine and modified affine Solomon-Stiffler codes. Projective Solomon-Stiffler codes are special cases of our modified affine Solomon-Stiffler codes. Several infinite families of $q$-ary Griesmer, optimal, almost optimal two-weight, three-weight, four-weight and five-weight linear codes are constructed as special cases of our construction. Weight distributions of these Griesmer, optimal or almost optimal codes are determined. Many optimal linear codes documented in Grassl’s list are re-constructed as (modified) affine Solomon-Stiffler codes.
@@ 1.1 Preliminaries
One of the main goal in the theory of error-correcting codes is to construct codes ${\bf C}\subset{\bf F}_{q}^{n}$ with large cardinalities and minimum distances. In general, there are some upper bounds on cardinalities or minimum distances of codes. Optimal codes attaining these bounds are particularly interesting, see [18]. A linear code over the finite field ${\bf F}_{q}$ with the length $n$, dimension $k$ and the minimum Hamming weight $d$ is denoted as a linear $[n,k,d]_{q}$ code. A linear code is called projective if its dual distance is at least three. Let $A_{i}({\bf C})$, $i=0,1,\ldots,n$, be the number weight $i$ codewords in the linear code ${\bf C}\subset{\bf F}_{q}^{n}$. Then it is obvious that there are at most $n-d({\bf C})+1$ nonzero weights $i\in\{d({\bf C}),d({\bf C})+1,\ldots,n\}$, such that $A_{i}({\bf C})\neq 0$. If $A_{i}({\bf C})\neq 0$ for exact $t$ indices $i_{1}<i_{2}<\cdots<i_{t}$, this linear code is called $t$-weight linear code. A simplex code is a one-weight code. When $t$ is small, constructions of few-weight linear codes and the determination of their exact weight distributions are important topic in coding theory.
@@ 1.2 Related works
Set $g_{q}(k,d)=\Sigma_{i=0}^{k-1}\lceil\frac{d}{q^{i}}\rceil$, for a linear $[n,k,d]_{q}$ code, the Griesmer defect is $g({\bf C})=n-g_{q}(k,d)$. A linear $[n,k,d]_{q}$ code attaining this bound, that is, the Griesmer defect zero code, is called a Griesmer code, see [30, 17]. In 1965, Solomon and Stiffler constructed infinite families of binary and $q$-ary Griesmer codes in [30]. In 1973, Baumert and McEliece proved that for fixed $q$ and $k$, there are Griesmer codes for any given fixed large minimum distance $d$, see [1]. Then many authors constructed Griesmer codes from various methods, see [13, 12, 11, 14, 2, 21, 19, 8, 16, 24, 17]. Optimal linear codes close to the Griesmer bound were also constructed in [27, 16, 17, 22]. However, codes in [20, 21] have Griesmer defects one or two.
In coding theory, it is extremely difficult to determine exact weight distributions of linear codes. For constructions of few-weight linear codes from Boolean functions and designs, and the determination of their exact weight distributions, we refer to [3, 4, 5, 6, 7, 31, 22, 23, 8, 16, 17, 32] and references therein.
@@
Suppose that there are $h$ dimension $u_{1},\ldots,u_{h}$ subspaces $S_{1},\ldots,S_{h}$ in ${\bf F}_{2}^{k}$, such that $S_{i}\bigcap S_{j}={\bf 0}$. Let $F$ be the set of disjoint nonzero column vectors in dimension $u_{1}$, $u_{2}$, …, $u_{h}$ subspaces $S_{1},\cdots,S_{h}$ in ${\bf F}_{2}^{k}$. Deleting these columns in the generator matrix of the binary simplex $[2^{k}-1,k,2^{k-1}]_{2}$ code, we obtain a binary linear $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code meeting the Griesmer bound, see [30]. Similarly, $q$-ary projective Solomon-Stiffler codes can be defined, we refer to [30]. In particular, when $h=1$, $u_{1}=k-1$, 1st order Reed-Muller code is a binary Solomon-Stiffler code associated with one positive integer $u_{1}=k-1$. However, weight distributions of general Solomon-Stiffler codes have not been determined completely.
@@
In [14, Theorem 3, 5], it was showed that projective binary or $q$-ary Griesmer codes are projective Solomon-Stiffler codes, when their minimum weights satisfy some condition. In the case of a binary Griesmer $[n,k,d]_{2}$ code, if $d\leq 2^{k-1}$, this code is binary Solomon-Stiffler or Belov code. For $q$-ary Griesmer codes, we refer conditions on $d$ to [14, Theorem 5]. Notice that Griesmer codes constructed in this paper are not projective. Then these Griesmer codes are not $q$-ary projective Solomon-Stiffler codes determined by Theorem 3 and Theorem 5 of [14].
@@
Several infinite families of $q$-ary linear Griesmer codes or optimal codes have been constructed in [27, 28, 26, 19, 24] as Gray images of two-weight codes over finite rings. Weight distributions of these Griesmer and optimal codes were determined in these two papers [27, 19]. In [27], Shi, Guan and Solé constructed optimal codes in [27, Theorem 6.1, 6.3] via Gray images of some two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$ and determined their exact weight distributions. The main contribution of [19] is the construction of Griemser codes in [19, Theorem 3.5], based a similar method. Similarly, Mesnager, Qian and Cao constructed optimal codes in [24, Theorem 5.4] via Gray images of codes over a finite chain ring. In [16], more general trace codes, which reach the Griesmer bound or are distance-optimal were constructed. In [26, Theorm 2] and [25, Proposition 4.3, 5)], some binary $[2^{k}-2^{u},k-1,2^{k-1}-2^{u-1}]_{2}$ codes were constructed. In a recent paper [15], punctured and shortening optimal codes were obtained from Griesmer cyclic codes.
1.3 Contributions and the organization of the paper
@@
2) In Sections 2 and 3, we show that these Griesmer codes and optimal codes constructed in [27, 19] and very special case codes in our construction have the same parameters and weight distributions. Optimal linear $q$-ary codes constructed in [24] have same parameters as very special codes in our construction. It is also indicated that Griesmer and optimal codes constructed in this paper are less restrictive than that constructed in [16]. In Section 7, we construct much more general distance-optimal binary codes than these constructed in [26, 25]. Our distance-optimal codes are codimension one or two subcodes of binary Solomon-Stiffler codes. Distance-optimal punctured codes of Griesmer codes are also discussed in Section 8. Our results in Section 7 and 8 are more general than main results of [15].
@@
4) Several infinite families of Griesmer or optimal $q$-ary $t$-weight linear codes, $q\leq 4$, $t\leq 5$, are constructed and their weight distributions are determined. Comparing with previous works on few-weight lineae codes in [3, 4, 5, 22, 23, 8, 32], codes in these infinite families are Griesmer, optimal or almost optimal. Moreover these codes are more general than codes constructed in previous papers, such as, [5, 27, 19, 28, 22, 24].
@@ 2 Affine Solomon-Stiffler codes
Let $s_{1}$ be the number of dimension $u_{1}$ subspaces, $s_{2}$ be the number of dimension $u_{s_{1}+1}$ subspaces,…, $s_{t}$ be the number of largest dimension $u_{s_{1}+s_{2}+\cdots+s_{t-1}+1}$ subspaces. Set $g=(s_{1}-1)+\cdots+(s_{t}-1)$.
Let $S$ be the set of all $q^{k}-1$ nonzero vectors in ${\bf F}_{q}^{k}$ deleting nonzero vectors in $S_{1}\bigcup\cdots\bigcup S_{h}$. Then
@@
From Lemma 1, column vectors in $S$ is a $k\times|S|$ matrix $G$ of the rank $k$. Let ${\bf C}$ be the code with the generator matrix $G$. For the convenience, we denote the above affine Solomon-Stiffler code by parameters $(k,{\bf u})=(k,u_{1},\ldots,u_{h})$. Then we have the following result.
Theorem 1 The linear code ${\bf C}$ is a linear $[q^{k}-1-\Sigma_{i=1}^{h}(q^{u_{i}}-1),k,d\geq(q-1)(q^{k-1}-\Sigma_{i=1}^{h}(q^{u_{i}-1})]_{q}$ code. This code has at most $2^{h}$ weights.
Proof. For any hyperplane $H\subset{\bf F}_{q}^{k}$, it is clear that $|H\bigcap S_{i}|$ is minimal if and only if $H\bigcap S_{i}$ is a hyperplane in $S_{i}$. Then $H-H\bigcap S$ is maximal, when each $H\bigcap S_{i}$ is a hyperplane in $S_{i}$, for $i=1,\ldots,h$. The weight of the nonzero codeword corresponding to $H$ is $|F_{2}^{k}-S|-|(H-H\bigcap S)|$. The first conclusion follows directly. Since each $|H\bigcap S_{i}|$ has two possible values, then $|H\bigcap S|$ has at most $2^{h}$ possible values. The second conclusion is proved.
@@
is not tight. For example, if $u_{1}=\cdots=u_{h}=1$, and $h$ is large. Many columns have to be deleted, a hyperplane cannot intersect each of such line at the zero point, see Section 6 below.
@@
$\Sigma_{i=0}^{k-1}\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=1}^{h}q^{u_{i}-1})}{q^{i}}\rceil$
, many Griesmer codes and optimal codes can be constructed, as in the following Theorem 2, Corollary 1 and 2.
@@
2) If $u_{1}=u_{2}=\cdots=u_{h}$, the Griesmer defect is upper bounded by $h-1$.
3) If $u_{1}=\cdots=u_{s_{1}}<u_{s_{1}+1}=\cdots=u_{s_{1}+s_{2}}<\cdots<u_{s_{1}+\cdots+s_{t-1}+1}=\cdots=u_{s_{1}+\cdots+s_{t}}$ and $u_{s_{1}+\cdots+s_{j}}\leq u_{s_{1}+\cdots+s_{j}+1}-2$. Then the Griesmer defect of the linear code ${\bf C}$ is upper bounded by $g$.
Proof. We calculate the Griesmer defect. Notice that $\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil\leq(q-1)\lceil\frac{(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil-(s_{2}-1)$, only when $i=u_{s_{1}+1}$. On the other hand, if $u_{s_{1}+s_{2}+\cdots+s_{j}+1}\geq u_{s_{1}+s_{2}+\cdots+s_{j}}+2$, this kind of decreasing happens only when $i=u_{s_{j}+1}$. Then the conclusion follows.
The calculation of the Griesmer defects reveals that, even affine Solomon-Stiffler codes introduced in this paper are not Griesmer as Solomon-Stiffler codes in [30]. However their Griesmer defects are small in many cases. Hence, many affine Solomon-Stiffler codes are close to the Griesmer bound, and are distance-optimal or distance-almost optimal, distance-near optimal, with respect to the Griesmer bound, when $h$ is small, see Table 1 below.
@@
Corollary 2 Suppose that $u_{1}\geq 2$. If only one $s_{j}$ is $2$ and other $s_{i}$’s are $1$. Moreover,
@@
Let $q$ be a prime power and $s$ be a positive integer satisfying $s\geq 3$. The affine Solomon-Stiffler code associated with parameters $(2s-2,s)$ is a linear Griesmer $[q^{2s-2}-q^{s-1},2(s-1),(q-1)(q^{2s-3}-q^{s-2}]_{q}$ code. The $q$-copy repetition code is a linear $[q^{2s-1}-q^{s},2(s-1),(q-1)(q^{2s-2}-q^{s-1})]_{q}$ code. This code is optimal when $s\geq q$ as argued in [24, Theorem 5.4]. Comparing these two codes, they have the same parameters. The construction of optimal codes in [24] is based on Griesmer codes over a finite chain ring. Our construction is simpler.
In [16, Theorem 1], similar codes with some strong restrictions on dimensions of subfields of ${\bf F}_{q^{k}}$ are studied. Only codes in the case when $u_{i}$ is not a factor of $u_{j}$, if $j>i$, or all $u_{i}$’s are the same, are studied. Comparing to [16], we need the key condition 1). However it is clear that Theorem 2 gives more general Griesmer and optimal codes than trace Griesmer and optimal codes given in [16].

### v3

@@ Abstract
In their fundamental paper published in 1965, G. Solomon and J. J. Stiffler invented infinite families of codes meeting the Griesmer bound. These codes are then called Solomon-Stiffler codes and have motivated various constructions of codes meeting or close the Griesmer bound. However weight distributions of Solomon-Stiffler codes have been only determined for very special cases.
In this paper, we give a geometric construction of affine and modified affine Solomon-Stiffler codes. Projective Solomon-Stiffler codes are special cases of our modified affine Solomon-Stiffler codes. Several infinite families of $q$-ary Griesmer, optimal, almost optimal two-weight, three-weight, four-weight and five-weight linear codes are constructed as special cases of our construction. Weight distributions of these Griesmer, optimal or almost optimal codes are determined explicitly. Many optimal linear codes documented in Grassl’s list are re-constructed as (modified) affine Solomon-Stiffler codes.
@@ 1.1 Preliminaries
One of the main goals in the theory of error-correcting codes is to construct codes ${\bf C}\subset{\bf F}_{q}^{n}$ with large cardinalities and minimum distances. In general, there are some upper bounds on cardinalities or minimum distances of codes. Optimal codes attaining these bounds are particularly interesting, see [22]. A linear code over the finite field ${\bf F}_{q}$ with the length $n$, dimension $k$ and the minimum Hamming weight $d$ is denoted as a linear $[n,k,d]_{q}$ code. A linear code is called projective if its dual distance is at least three. Let $A_{i}({\bf C})$, $i=0,1,\ldots,n$, be the number weight $i$ codewords in the linear code ${\bf C}\subset{\bf F}_{q}^{n}$. Then it is obvious that there are at most $n-d({\bf C})+1$ nonzero weights $i\in\{d({\bf C}),d({\bf C})+1,\ldots,n\}$, such that $A_{i}({\bf C})\neq 0$. If $A_{i}({\bf C})\neq 0$ for exact $t$ indices $i_{1}<i_{2}<\cdots<i_{t}$, this linear code is called $t$-weight linear code. A simplex code is an one-weight code. When $t$ is small, constructions of few-weight linear codes and the determination of their exact weight distributions are important topic in coding theory.
@@ 1.2.1 Griesmer codes and Solomon-Stiffler codes
Set $g_{q}(k,d)=\Sigma_{i=0}^{k-1}\lceil\frac{d}{q^{i}}\rceil$, for a linear $[n,k,d]_{q}$ code, the Griesmer defect is $g({\bf C})=n-g_{q}(k,d)$. A linear $[n,k,d]_{q}$ code attaining this bound, that is, the Griesmer defect zero code, is called a Griesmer code, see [39, 21]. In 1965, Solomon and Stiffler constructed infinite families of binary and $q$-ary Griesmer codes in [39]. In 1973, Baumert and McEliece proved that for fixed $q$ and $k$, there are Griesmer codes for any given fixed large minimum distance $d$, see [1]. Then many authors constructed Griesmer codes from various methods, see [17, 16, 15, 18, 5, 26, 23, 11, 20, 29, 21, 10]. Optimal linear codes close to the Griesmer bound were also constructed in [36, 20, 21, 27]. However, codes in [24, 26] have Griesmer defects one or two.
In coding theory, it is extremely difficult to determine exact weight distributions of linear codes. For constructions of few-weight linear codes from Boolean functions and designs, and the determination of their exact weight distributions, we refer to [6, 8, 9, 10, 40, 27, 28, 11, 20, 21, 41] and references therein.
@@
Suppose that there are $h$ dimension $u_{1},\ldots,u_{h}$ subspaces $S_{1},\ldots,S_{h}$ in ${\bf F}_{2}^{k}$, such that $S_{i}\bigcap S_{j}={\bf 0}$. Let $F$ be the set of disjoint nonzero column vectors in dimension $u_{1}$, $u_{2}$, …, $u_{h}$ subspaces $S_{1},\cdots,S_{h}$ in ${\bf F}_{2}^{k}$. Deleting these columns in the generator matrix of the binary simplex $[2^{k}-1,k,2^{k-1}]_{2}$ code, we obtain a binary linear $[2^{k}-1-\Sigma_{i=1}^{h}(2^{u_{i}}-1),k,2^{k-1}-\Sigma_{i=1}^{h}2^{u_{i}-1}]_{2}$ code meeting the Griesmer bound, see [39]. Similarly, $q$-ary projective Solomon-Stiffler codes can be defined, we refer to [39]. In particular, when $h=1$, $u_{1}=k-1$, 1st order Reed-Muller code is a binary Solomon-Stiffler code associated with one positive integer $u_{1}=k-1$. However, weight distributions of general Solomon-Stiffler codes have not been determined completely. Binary Solomon-Stiffler codes were also used to construct optimal binary linear complementary pairs in [14].
@@
In [18, Theorems 3, 5], it was showed that projective binary or $q$-ary Griesmer codes are projective Solomon-Stiffler codes, when their minimum weights satisfy some condition. In the case of a binary Griesmer $[n,k,d]_{2}$ code, if $d\leq 2^{k-1}$, this code is binary Solomon-Stiffler or Belov code. For $q$-ary Griesmer codes, we refer conditions on $d$ to [18, Theorem 5]. Notice that Griesmer codes constructed in this paper are not projective. Then these Griesmer codes are not $q$-ary projective Solomon-Stiffler codes determined by Theorem 3 and Theorem 5 of [18].
1.2.2 Optimal codes from codes over finite rings
@@
Several infinite families of $q$-ary linear Griesmer codes or optimal codes have been constructed in [36, 37, 35, 23, 29] as Gray images of two-weight codes over finite rings. Weight distributions of these Griesmer and optimal codes were determined in these two papers [36, 23]. In [36], Shi, Guan and Solé constructed optimal codes in [36, Theorem 6.1, 6.3] via Gray images of some two-weight codes over the finite ring ${\bf F}_{p}+u{\bf F}_{p}$ and determined their exact weight distributions. The main contribution of [23] is the construction of Griemser codes in [23, Theorem 3.5], based on a similar method. Similarly, Mesnager, Qian and Cao constructed optimal codes in [29, Theorem 5.4] via Gray images of codes over a finite chain ring. In [20], more general trace codes, which reach the Griesmer bound or are distance-optimal were constructed. In [35, Theorm 2] and [30, Proposition 4.3, 5)], some binary $[2^{k}-2^{u},k-1,2^{k-1}-2^{u-1}]_{2}$ codes were constructed. In a recent paper [19], punctured and shortening optimal codes were obtained from Griesmer cyclic codes.
1.2.3 Grassmann codes and related codes
Projective and affine Grassmann codes and their duals were proposed and studied in [31, 2, 3] by D. Y. Nogin, P. Beelen, S. Ghorpade and T. Høholdt. Weight spectrum were calculated for some special Grassmann codes, see [25, 32]. Fast decoding algorithms of Grassmann codes were studied in [4, 33].
Let $m$ be a positive integer and $l<m$ be another positive integer, then linear affine Grassmann $[q^{l(m-l)},\displaystyle{m\choose l},q^{l(m-l)}\prod_{i=1}^{l}(1-\frac{1}{q^{i}})]_{q}$ codes $C(l,m)$ were constructed. It is interesting to observe that when $l$ is small, for example, $l=2$ and $m=4$, these Grassmann $[q^{4},6,(q-1)(q^{3}-q)]_{q}$ codes have relative small Grisemer defects at most $q^{2}-q$. Therefore, these affine codes are optimal or close to optimal. For example, when $q=2$, an affine Grassmann $[16,6,6]_{2}$ code is obtained. This is an optimal binary code, see [12]. When $m=3$, an affine $[81,6,48]_{3}$ code is obtained. The optimal ternary code in this case is a linear $[81,6,51]_{3}$ code, see [12]. On the other hand, affine Solomon-Stiffler $[q^{4}-q^{2},4,(q-1)(q^{3}-q)]_{q}$ code is constructed in Section 2. It seems interesting to find some possible relations between affine Grassmann codes and affine Solomon-Stiffler codes. There have been many papers on codes related to Grassmannians, see [34] and references therein.
1.3 Our contributions
@@
2) In Sections 2 and 3, we show that these Griesmer codes and optimal codes constructed in [36, 23] and very special case codes in our construction have the same parameters and weight distributions. Optimal linear $q$-ary codes constructed in [29] have same parameters as very special codes in our construction. It is also indicated that Griesmer and optimal codes constructed in this paper are less restrictive than those constructed in [20]. In Section 7, we construct much more general distance-optimal binary codes than those constructed in [35, 30]. Our distance-optimal codes are codimension one or two subcodes of binary Solomon-Stiffler codes. Distance-optimal punctured codes of Griesmer codes are also discussed in Section 8. Our results in Section 7 and 8 are more general than main results of [19].
@@
4) Several infinite families of Griesmer or optimal $q$-ary $t$-weight linear codes, $q\leq 4$, $t\leq 5$, are constructed and their weight distributions are determined. Comparing with previous works on few-weight linear codes in [6, 8, 9, 27, 28, 11, 41], codes in these infinite families are Griesmer, optimal or almost optimal. Moreover these codes are more general than codes constructed in previous papers, such as, [9, 36, 23, 37, 27, 29].
@@ 2 Affine Solomon-Stiffler codes
Let $s_{1}$ be the number of dimension $u_{1}$ subspaces, $s_{2}$ be the number of dimension $u_{s_{1}+1}$ subspaces,…, $s_{t}$ be the number of largest dimension $u_{s_{1}+s_{2}+\cdots+s_{t-1}+1}$ subspaces, where $s_{1}+s_{2}+\cdots+s_{t}\leq h$. Set $g=(s_{1}-1)+\cdots+(s_{t}-1)$. In the case of $q=2$, an important difference to Solomon-Stiffler codes is that the condition $s_{i}\leq 1$ was imposed in [39].
Let ${\bf g}_{1},\ldots,{\bf g}_{q^{k}-1}$ be all nonzero vectors of ${\bf F}_{q}^{k}$. Let $S$ be the set of all vectors among these $q^{k}-1$ nonzero vectors in ${\bf F}_{q}^{k}$ by deleting these nonzero vectors in $S_{1}\bigcup\cdots\bigcup S_{h}$. Then
@@
From Lemma 1, the column vectors in $S$ form a $k\times|S|$ matrix $G$ of the rank $k$. Let ${\bf C}$ be the code with the generator matrix $G$. For the convenience, we denote the above affine Solomon-Stiffler code by parameters $(k,{\bf u})=(k,u_{1},\ldots,u_{h})$. Then we have the following result.
Theorem 1 The linear code ${\bf C}$ is a linear $[q^{k}-1-\Sigma_{i=1}^{h}(q^{u_{i}}-1),k,d\geq(q-1)(q^{k-1}-\Sigma_{i=1}^{h}q^{u_{i}-1})]_{q}$ code. This code has at most $2^{h}$ weights.
Proof. For any hyperplane $H\subset{\bf F}_{q}^{k}$, it is clear that $|H\bigcap S_{i}|$ is minimal if and only if $H\bigcap S_{i}$ is a hyperplane in $S_{i}$. Then $H-H\bigcap S$ is maximal, when each $H\bigcap S_{i}$ is a hyperplane in $S_{i}$, for $i=1,\ldots,h$. The weight of the nonzero codeword corresponding to $H$ is $|S|-|(H-H\bigcap S)|$. The first conclusion follows directly. Since each $|H\bigcap S_{i}|$ has two possible values, then $|H\bigcap S|$ has at most $2^{h}$ possible values. The second conclusion is proved.
@@
is not tight. For example, if $u_{1}=\cdots=u_{h}=1$, and $h$ is large, many columns have to be deleted, a hyperplane cannot intersect each of such line at the zero point, see Section 6 below.
@@
$\Sigma_{i=0}^{k-1}\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=1}^{h}q^{u_{i}-1})}{q^{i}}\rceil,$
many Griesmer codes and optimal codes can be constructed, as in the following Theorem 2, Corollary 1 and 2.
@@
2) If $u_{1}=u_{2}=\cdots=u_{h}$ and $h\leq q$, the Griesmer defect is upper bounded by $h-1$.
3) If $u_{1}=\cdots=u_{s_{1}}<u_{s_{1}+1}=\cdots=u_{s_{1}+s_{2}}<\cdots<u_{s_{1}+\cdots+s_{t-1}+1}=\cdots=u_{s_{1}+\cdots+s_{t}}$ and $u_{s_{1}+\cdots+s_{j}}\leq u_{s_{1}+\cdots+s_{j}+1}-2$ for $j=1,2,\ldots,t-1$. Then the Griesmer defect of the linear code ${\bf C}$ is upper bounded by $g$.
Proof. 1) We observe that $\lceil\frac{(q-1)(q^{k-1}-q^{u_{1}-1}-q^{u_{2}-1}-\cdots-q^{u_{h}-1})}{q^{u_{1}}}\rceil=(q-1)q^{k-1-u_{1}}-\\ \lfloor\frac{(q-1)(q^{u_{1}-1}+q^{u_{2}-1}+\cdots+q^{u_{h}-1})}{q^{u_{1}}}\rfloor=(q-1)q^{k-1-u_{1}}$, since $(q-1)(q^{u_{1}-1}+\cdots+q^{u_{h}-1})<q^{u_{1}}$, from the condition $u_{1}<u_{2}<\cdots<u_{h}$. Then the right side of the Griesmer bound is $\Sigma_{i=0}^{k-1}\lceil\frac{(q-1)(q^{k-1}-q^{u_{1}-1}-\cdots-q^{u_{h}-1})}{q^{i}}\rceil=(q-1)(\frac{q^{k}-1}{q-1}-\Sigma_{j=1}^{h}\frac{q^{u_{j}}-1}{q-1})=(q^{k}-1)-\Sigma_{j=1}^{h}(q^{u_{j}}-1)$.
2) We observe that $\lceil\frac{(q-1)(q^{k-1}-hq^{u_{1}-1})}{q^{u_{1}}}\rceil=(q-1)q^{k-1-u_{1}}-\lfloor\frac{h(q-1)(q^{u_{1}-1})}{q^{u_{1}}}\rfloor=(q-1)q^{k-1-u_{1}}-\lfloor\frac{h(q-1)}{q}\rfloor\leq(q-1)q^{k-1-u_{1}}-h+1$. On the other hand, the term $\lceil\frac{(q-1)(q^{k-1}-hq^{u_{1}-1})}{q^{u_{1}+1}}\rceil=(q-1)q^{k-u_{1}-2}-\lfloor\frac{h(q-1)}{q^{2}}\rfloor=(q-1)q^{k-2-u_{1}}$, from the condition $h\leq q$. Then the conclusion follows directly.
3) We calculate the Griesmer defect. Notice that $\lceil\frac{(q-1)(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil\\ \leq(q-1)\lceil\frac{(q^{k-1}-\Sigma_{i=s_{1}+1}^{s_{1}+s_{2}}q^{u_{i}-1})}{q^{i}}\rceil-(s_{2}-1)$, only when $i=u_{s_{1}+1}$. On the other hand, if $u_{s_{1}+s_{2}+\cdots+s_{j}+1}\geq u_{s_{1}+s_{2}+\cdots+s_{j}}+2$, this kind of decreasing happens only when $i=u_{s_{j}+1}$. Then the conclusion follows.
The calculation of the Griesmer defects reveals that, affine Solomon-Stiffler codes introduced in this paper are not Griesmer as Solomon-Stiffler codes in [39]. However their Griesmer defects are small in many cases. Hence, many affine Solomon-Stiffler codes are close to the Griesmer bound, and are distance-optimal or distance-almost optimal, distance-near optimal, with respect to the Griesmer bound, when $h$ is small, see Table 1 below.
@@
Corollary 2 Suppose that $u_{1}\geq 2$ and only one $s_{j}$ is $2$ and other $s_{i}$’s are $1$. Moreover,
@@
Let $q$ be a prime power and $s$ be a positive integer satisfying $s\geq 3$. The affine Solomon-Stiffler code associated with parameters $(2s-2,s)$ is a linear Griesmer $[q^{2s-2}-q^{s-1},2(s-1),(q-1)(q^{2s-3}-q^{s-2}]_{q}$ code. The $q$-copy repetition code is a linear $[q^{2s-1}-q^{s},2(s-1),(q-1)(q^{2s-2}-q^{s-1})]_{q}$ code. This code is optimal when $s\geq q$ as argued in [29, Theorem 5.4]. Two optimal linear codes have the same parameters. The construction of optimal codes in [29] is based on Griesmer codes over a finite chain ring. Our construction is simpler.
@@
In [20, Theorem 1], similar codes with some strong restrictions on dimensions of subfields of ${\bf F}_{q^{k}}$ are studied. Only codes in the case when $u_{i}$ is not a factor of $u_{j}$, if $j>i$, or all $u_{i}$’s are the same, are studied. Comparing with [20], we also need the key condition 1). However it is clear that Theorem 2 gives more general Griesmer and optimal codes than trace Griesmer and optimal codes given in [20]. It also should be indicated that distance-optimal codes constructed in [7, Theorem 9] have the same parameters as very special codes of our construction. The binary Griesmer codes constructed in [7, Theorem 7] are actually classical binary Solomon-Stiffler codes.